### v1

# MegaSynth: Scaling Up 3D Scene Reconstruction with Synthesized Data

Abstract
We propose scaling up 3D scene reconstruction by training with synthesized data. At the core of our work is MegaSynth, a procedurally generated 3D dataset comprising 700K scenes—over 50 times larger than the prior real dataset DL3DV—dramatically scaling the training data. To enable scalable data generation, our key idea is eliminating semantic information, removing the need to model complex semantic priors such as object affordances and scene composition. Instead, we model scenes with basic spatial structures and geometry primitives, offering scalability. Besides, we control data complexity to facilitate training while loosely aligning it with real-world data distribution to benefit real-world generalization. We explore training LRMs with both MegaSynth and available real data. Experiment results show that joint training or pre-training with MegaSynth improves reconstruction quality by 1.2 to 1.8 dB PSNR across diverse image domains. Moreover, models trained solely on MegaSynth perform comparably to those trained on real data, underscoring the low-level nature of 3D reconstruction. Additionally, we provide an in-depth analysis of MegaSynth’s properties for enhancing model capability, training stability, and generalization.
Figure 1: We introduce MegaSynth, a non-semantic synthesized dataset for training LRMs. MegaSynth benefits from its scalability and controllability, enabling us to generate 700K scenes in 3 days. We train LRMs with both the large-scale MegaSynth data and small-scale real data, improving LRMs for reconstructing wide-coverage scenes from dense-view images.
1 Introduction
The scaling law has shifted the focus of contemporary AI research toward large foundation models, which are built with scalable neural network architectures [67, 25] and trained on vast datasets [58, 5]. Following the scaling recipe seen in NLP and 2D vision [1, 3, 4], the Large Reconstruction Model (LRM) has been introduced to learn general 3D reconstruction priors [27]. For object-level reconstruction, LRMs have shown impressive reconstruction quality using either single-view or sparse-view inputs [27, 69, 85, 33], enabling a range of applications [38, 87].
Despite progress, enhancing LRM for reconstructing wide-coverage scenes remains challenging due to two key limitations of training data. First, scene-level datasets are significantly smaller in scale compared to object-level counterparts. For instance, Objaverse [17] contains 800K shape instances, whereas the largest clean scene dataset, DL3DV, includes just 10K scenes. Collecting more intentionally captured scene data is labor-intensive and difficult to scale. Second, existing scene-level datasets suffer from a suboptimal data distribution. They are often limited by insufficient scene diversity [15], small camera motions [88, 47], noisy content [66], and inaccurate annotations [39]. However, given the inherent complexity of 3D scenes, effective training requires clean and diverse data, especially multi-view images captured by widely spaced cameras with precise camera annotations [48].
In this work, we propose scaling up training data for scene-level reconstruction by using synthesized data. Our key idea is to eliminate the reliance on semantic information in data generation, by constructing scenes with non-semantic shape primitives arranged within basic spatial structures. This approach is motivated by our insight that scene semantics play a minimal role in multi-view reconstruction, as evidenced by the success of traditional non-semantic methods such as COLMAP [57], MVS [59], NeRF [48], and the emerging non-semantic properties of recent feed-forward models [29, 9, 76]. Unlike prior scene generation methods, which aim to replicate real-world scene distributions [54, 63, 16, 72, 53] and are thus constrained by the complexity of modeling semantics, such as object affordances, our approach bypasses these challenges. This simplification enables highly scalable and efficient data generation.
Beyond scalability, synthesized data offers controllability. We control data complexity to facilitate training while loosely aligning it with real-world data distribution to benefit real-world generalization. Through heuristic methods, we regulate key factors, such as geometric complexity, camera pose distribution, materials, and lighting, for creating diverse scenes. Additionally, synthesized data provides precise metadata, such as camera and geometry information, further ensuring improved training stability and effectiveness.
We generate the MegaSynth dataset, comprising 700K scenes. MegaSynth is over 50 times larger than the real dataset DL3DV and significantly scales up training data for LRMs. We utilize MegaSynth to train feed-forward LRMs [85, 89] jointly with DL3DV. Our experiments show a 1.2 to 1.8 dB PSNR gain across diverse test datasets and image resolutions, highlighting the synergy between synthesized and real data. Synthesized data excels in scale and provides rich metadata, such as geometry supervision, enabling models to develop a general geometric understanding beyond rendering supervision. Meanwhile, small-scale real data further sharpens the model. Notably, models trained solely on MegaSynth, regardless of the number of input views, perform comparably to those trained on real data. This result supports our hypothesis that 3D reconstruction is predominantly a low-level vision task, with minimal reliance on semantics.
2 Related Work
Scene-level 3D Reconstruction. Reconstructing scenes has been a long-standing challenge in 3D computer vision. Traditional Structure-from-Motion (SfM) and Multi-view Stereo (MVS) methods, as well as their neural counterparts, adopt a bottom-up approach [59, 57, 62, 21, 20, 31, 61, 68]. For instance, COLMAP [57] builds from low-level visual cues to more structured geometry through keypoint detection, matching, camera reconstruction, and bundle adjustment.
Learning-based methods encompass both 3D neural scene representations and feed-forward prediction models. Researchers have explored the distinct properties of explicit [71, 49], implicit [60, 8, 48, 42], and hybrid 3D representations [35, 23, 28, 32] to enhance reconstruction quality, typically optimizing the 3D representation for each scene to demonstrate capability. Meanwhile, generalizable reconstruction models have been developed, where neural networks predict 3D representation attributes in a feed-forward manner. Some approaches follow a traditional bottom-up paradigm, leveraging inductive biases from MVS [10, 84, 70], cost volume [12, 13], correspondence cues [11], and epipolar geometry [9, 75, 19]. In contrast, recent work proposes top-down frameworks [29, 69, 71, 37] that infer geometry directly and better harness the power of large models. However, some of these works rely on pairwise computations [71], which limits a global understanding of inputs. Our work, in contrast, leverages recent global-aware methods [85, 89] and focuses on scaling up training data to advance dense-view reconstruction.
Large Reconstruction Model (LRM). LRMs have been introduced to scale up generalizable 3D reconstruction methods [27], employing scalable network architectures and training on large datasets to learn generic reconstruction priors. Typically, LRMs use Transformers [67, 27, 29] or U-Nets [55, 64] as model backbones, encoding 2D image inputs into 3D representations, e.g., Triplane [27, 69] and mesh [73, 77], enabling high-quality object reconstruction. The following research has focused on enhancing object reconstruction by incorporating generative priors [78, 86] and designing more scalable training frameworks [76, 30, 24]. Additionally, novel 3D representations, such as 3D Gaussians [35], have extended LRMs to scene-level 3D reconstruction [85, 89]. However, reconstructing wide-coverage scenes remains challenging due to limited data. To address this, we propose a scalable data generation method considering the non-semantic property of multi-view reconstruction.
Training with Synthesized Data. Leveraging synthesized data for training is essential when available data is insufficient or biased. Synthesized data has been widely applied across fields such as Robotics [63], Natural Language Processing [2], Computer Vision [46], and AI for Science [65]. For example, recent depth estimation methods utilize synthesized data’s accurate ground truth to enhance performance on fine structures [82, 7]. A relevant topic to our work is 3D scene generation, where generated data supports training 3D reconstruction models [50, 16, 52, 81, 44, 51, 26]. However, these methods focus on generating realistic scenes, necessitating semantic accuracy (e.g., object affordance and relationships), which constrains scalability due to the complex procedural rules required for accuracy and diversity. While some recent methods attempt to address this limitation with language models [83], these models often lack spatial awareness and are slow in inference. In contrast, we show that semantics are not essential for multi-view reconstruction, allowing us to create a data generation pipeline free from semantic constraints and capable of generating virtually unlimited training data. Previously, non-semantic shape primitives have been used for various object reconstruction and appearance acquisition tasks [79, 80, 40, 56, 41]. Recently, LRM-Zero [76] has used primitive-based methods to generate large-scale data to train large reconstruction models, but it is limited to the object level. We focus on more challenging scene-level data synthesis, incorporating control of lighting, object composition, and camera poses for reconstructing wide-coverage scenes from dense-view images. We also present a mixed training framework to leverage the synergy between synthesized and real data. DUST3R [71] employs a pre-trained encoder from CroCo [74], which incorporates synthesized data, but its pre-training is limited to 2D image representation learning without directly learning 3D priors. In contrast, our pre-training approach directly targets 3D scenes, enhancing our model’s geometric and texture understanding. We also enable joint training with both synthesized and real data.
3 Task and Preliminary
Our goal is to reconstruct wide-coverage scenes in a feed-forward manner. Given a set of dense-view images $\{I_{i}\mid i=1,...,n\}$ with known camera information, the model predicts the attributes of 3D representations. By default, we use $\mathbf{n=32}$ views in our experiments to handle the high complexity of scenes, in contrast to previous sparse-view methods that rely on only 4 to 8 views [73, 69].
This paper primarily experiments with GS-LRM [85] and Long-LRM [89], chosen for their strong reconstruction quality. Both methods predict pixel-aligned 3D Gaussians from posed images with similar model architectures but different backbones: GS-LRM and Long-LRM employ transformer-based and Mamba-based [22, 43] backbones, respectively.
Given the input views, the models first patchify each image using non-overlapping convolutions, encoding them into feature tokens $\{T_{i}\mid i=1,...,n\}$ as in ViT [18]. The feature tokens from all images are flattened and concatenated into a feature set, $\mathbf{F}$, which is later processed by the model $\mathcal{M}$. Finally, an MLP decodes Gaussian parameters $\mathbf{G}$ to represent the scene. The process is formulated as follows:
$\{T_{1},\dots,T_{n}\}$ =$=\{\mathrm{Conv}(I_{1}),\dots,\mathrm{Conv}(I_{n})\},$ (1)
$\mathbf{F}$ $=[\mathrm{Flatten}(T_{1}),\dots,\mathrm{Flatten}(T_{n})],$ (2)
$\bar{\mathbf{F}}$ $=\mathcal{M}(\mathbf{F}),$ (3)
$\mathbf{G}$ $=\mathrm{MLP}(\bar{\mathbf{F}}),$ (4)
where [,,][$\cdot$,...$\cdot$][ ⋅ , … , ⋅ ] denotes concatenation, and $\bar{\mathbf{F}}$ represents the updated feature tokens produced by the backbone.
In the next section, we introduce our approach to synthesize data for training these models.
4 Synthesizing the MegaSynth Dataset
In this section, we first give an overview of our data synthesis method and then dive deeper to introduce how we control complexity, diversity, and alignment with real data.
Figure 2: MegaSynth generation pipeline. We first generate the scene floor plan, where each 3D box represents a shape and different colors represent different object types. We compose shape primitives into objects with geometry augmentations, where these objects further compose the scene. We randomize the texture and lighting, and generate random cameras for rendering.
4.1 Overview
We synthesize MegaSynth using a procedural generation method, as illustrated in Fig. 2. The process involves: i) generating a scene floor plan, including the scene size and object instance box locations, ii) instantiating object geometries with random textures, and iii) randomizing the lighting. During this process, we eliminate high-level scene semantics, e.g., category-related object affordance. We only keep the low-level structural and geometric features of scenes.
4.2 Scene Floor Plan
Without loss of generality, we plan the scene as a cube box and populate it with objects represented by 3D bounding boxes. We randomize the 3D aspect ratio and size of scenes. We design multiple object box categories to simulate real-world scene geometry structures (visualized as boxes in Fig. 2 with different colors). For example, large object boxes tend to be placed near the ground while small object boxes have more flexible placement options. We parameterize the size, location, and number of each object type, specifying each parameter as a range. This allows us to introduce randomness to improve diversity. Further details of the object box categories and their attribute sampling ranges are provided in the Appendix A.
4.3 Geometry and Texture
The scene floor plan constructed above divides the room space into basic units of object boxes. We then synthesize geometry and assign textures for each geometric shape.
Geometry of general objects. For each object box, we generate geometry by combining non-semantic shape primitives [76, 79], including cubes, spheres, cylinders, and cones. These primitives incorporate diverse geometry patterns, such as flat and curved surfaces, straight and curved lines, and sharp edges. Composing these shapes further increases geometric and topological complexity. Additionally, we apply random height-field augmentations [79] to the primitives, producing surfaces with both concave and convex details.
Different object categories (defined in Sec. 4.2) utilize varying numbers of shape primitives; for instance, large objects are typically composed of more primitives than small ones, loosely reflecting the complexity distribution of real-world objects. The geometry is instantiated in a canonical space, then rescaled and translated to fit the object box.
Geometry for increasing complexity. To enhance diversity and alignment with real data, we incorporate two additional types of geometry. First, we add thin structures, such as wireframes of shape primitives, enabling the reconstruction of fine-grained geometries. To further increase diversity, we randomly place solid primitives intersecting with these wireframes. Second, we introduce axis-aligned geometries, such as thin sticks and flat surfaces, to reflect real-world geometry distributions under the Manhattan assumption [14].
Texture. Each shape primitive is assigned a random texture, including a basic color map along with normal, material, and roughness maps. We increase the probability of sampling specular and glass materials, ensuring a closer match to real-world appearances.
4.4 Lighting
Real-world images often feature complex lighting conditions. Thus, we design three lighting conditions and randomly compose them to improve the diversity and complexity. Each lighting uses a randomly sampled color and intensity.
Ambient light. We use the uniformly distributed ambient lighting with a unit brightness by default. The ambient lighting provides consistent illumination across a scene, helping to reveal scene details and stabilizing training.
Sunlight. Adding sunlight simulates true-to-life lighting effects, making the scene more complex with a higher intensity and casting shadows. We set the sunlight outside of the scene box. To enable the sunlight effect within the scene, we create windows on the walls with random sizes, under the regions that the sunlight covers. To further improve the complexity and diversity, we randomly add window bars implemented as the wireframes and window glasses.
Luminous objects and light bulbs. We randomly turn some objects and axis-aligned sticks in the scene as lights to simulate real-world lighting. We also randomly place light bulbs in the scene to increase diversity. The intensity of object light can be sampled as large values to simulate lighting in dark environments.
5 Learning 3D Reconstruction on MegaSynth
In this section, we discuss how we utilize our synthesized MegaSynth (Sec. 4) to train a feed-forward reconstruction model (i.e., the LRM-based model illustrated in Sec. 3). To reach the goal, we first construct the training data by carefully sampling the camera distribution and rendering the images (Sec. 5.1). We then train our model with a mixed-data training strategy (Sec. 5.2) with rendering loss and geometry loss (Sec. 5.3). The details of the training process can be found in the Appendix B.
5.1 Training Data Preparation on MegaSynth
To get training data, we render input views and target supervisions from the synthesized MegaSynth scenes. We sample cameras and then render RGB and depth images accordingly. We do not distinguish the input views and target views, i.e., they will be used interchangeably during training.
The main challenge of this data creation pipeline is the camera pose sampling. We empirically found that a careful design of camera sampling distribution can largely improve learning efficiency, model generalization, and training stability. We next detail our camera sampling process.
Basic rules. The cameras are sampled to keep a minimal distance from any objects in the scene, preventing the camera from losing context and avoiding the near-clipping issue. We randomly sample the field-of-view (FoV) of cameras, due to the diversity of lenses used in real-world image capture.
Better scene coverage. We heuristically split the scene into the inner and outer spaces, based on the distance to the scene center. The cameras sampled in outer space always look at the scene center, ensuring better view coverage. Meanwhile, the cameras in the inner space are encouraged to have more diverse poses, e.g. the orientations are randomly sampled within pre-defined ranges, increasing the diversity and matching real-world camera pose distribution.
Constrained camera baseline. The randomly sampled cameras in the outer part of the scene tend to have large baselines. To improve diversity, we choose to sample more scenes and cameras with slightly smaller baselines, aligning with real-world camera distribution. Thus, instead of sampling camera position in all free space, we first sample a distance range and then sample the camera within the constrained space.
5.2 Mixed Data Training
In training, we leverage distinct advantages from both the synthesized MegaSynth renderings and the real-world dataset (e.g., DL3DV). The synthesized data, with its diversity and scale, provide a foundation for models to learn general reconstruction priors of geometry, texture, and lighting. Moreover, easy access to accurate metadata (e.g., depth images and noise-free camera information) enhances geometric understanding and stabilizes training.
Meanwhile, real-world data offers authenticity that is hard to synthesize yet crucial for model robustness. For instance, it captures real-world imperfections like sensor noise and lighting artifacts, enhancing the model’s robustness for real-world deployment. Additionally, its realistic semantics better align the model with real-world scene distributions.
We find these datasets to be complementary. Our experiments investigate two training strategies to leverage their synergy: (1) pre-training on the large-scale MegaSynth dataset followed by fine-tuning on a smaller real-world dataset; and (2) joint training on both datasets simultaneously. These approaches balance scalability and authenticity.
5.3 Rendering and Geometry Losses
We follow the standard method for training large reconstruction models using photometric image rendering losses:
[…]
where $\lambda$ is the weight for balancing the perceptual loss [34], $I_{i}$ is ground-truth target image, and $\hat{I}_{i}$ is image rendered from predicted 3D Gaussians under target camera poses.
Our synthesized data naturally provides accurate geometry information, which is utilized to supervise the geometry of the 3D Gaussians predicted by the LRM models. In detail, both GS-LRM and Long-LRM (described in Sec. 3) predict pixel-aligned 3D Gaussians, where each Gaussian corresponds to a pixel in the input view. We supervise the center location of the predicted 3D Gaussians using the ground-truth geometry information. It is formulated as
[…]
6 Experiments
Figure 3: Reconstruction visualization on the in-domain DL3DV data. The results are from Long-LRM at resolution 256. We present both indoor and outdoor results in the first and second rows, respectively. With our MegaSynth (denoted as ‘w. MegaSynth’), the model performs better on thin structures (e.g., bottom left), complicated lighting (e.g., top middle), and cluttered scenes (e.g., top right).
Inf. Time In-Domain Out-of-Domain (Zero-shot Generalization)
DL3DV Hypersim MipNeRF360 & TT
Model Training Dataset PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Resolution 128, 32 input views
3DGS [35] N.A. (Per-scene Optimization) 5.2 min 24.27 0.817 0.166 20.67 0.672 0.293 16.46 0.458 0.405
Long-LRM [89] DL3DV 0.12 sec 24.18 0.812 0.173 23.41 0.790 0.210 19.68 0.569 0.312
Long-LRM (ours) DL3DV + MegaSynth 25.44 0.853 0.136 25.01 0.836 0.164 20.86 0.652 0.249
GS-LRM [85] DL3DV 0.11 sec 24.60 0.824 0.161 23.89 0.806 0.195 19.93 0.601 0.289
GS-LRM (ours) DL3DV + MegaSynth 25.75 0.859 0.130 25.46 0.846 0.154 21.19 0.672 0.235
Resolution 256, 32 input views
3DGS [35] N.A. (Per-scene Optimization) 6.4 min 23.26 0.778 0.206 21.75 0.690 0.294 16.06 0.436 0.421
Long-LRM [89] DL3DV 0.35 sec 23.71 0.779 0.236 22.51 0.767 0.267 18.61 0.465 0.421
Long-LRM (ours) DL3DV + MegaSynth 25.14 0.828 0.186 24.26 0.817 0.210 19.84 0.555 0.339
Table 1: Evaluation results against baseline methods. We report results resolution of 128 and 256. For resolution 256, we only report results of Long-LRM as transformer-based GS-LRM is too slow. Our models are pre-trained on MegaSynth and then tuned on DL3DV.
In this section, we describe the experimental setting and present evaluation results. Due to the space limit, implementation and training details are in the Appendix B.
Figure 4: Reconstruction visualization on the out-of-domain data. The results are from Long-LRM at resolution 256. We include results for both Hypersim and MipNeRF360 are presented in the first and second rows, respectively.
MegaSynth-only Training Real Data Tuning
Data $\mathcal{L}^{S}_{\text{loc}}$ Scale (Trained w. only MegaSynth) (Using DL3DV)
Control Up Fail. Iter. PSNR SSIM LPIPS Fail. Iter. PSNR SSIM LPIPS
(0) 70k 17.18 0.519 0.445 80k 18.44 0.577 0.418
(1) 45k 18.71 0.601 0.384 57k 21.87 0.738 0.266
(2) - 20.72 0.691 0.300 - 25.12 0.835 0.171
(3) - 21.07 0.698 0.292 - 25.46 0.846 0.154
Table 2: Ablation study on data control, property and quantity. Results are reported on the Hypersim dataset with resolution 128. We also report the number of iterations before the job fails. Please see Appendix for data control details for experiment (0). The default data is composed of 100K examples, and the scaled one contains 700K examples.
6.1 Datasets
Besides our MegaSynth, we use three datasets in our paper, where DL3DV is the only one we take into training, i.e., others are evaluation-only.
DL3DV [45] [^1] We refer to the DL3DV-10K dataset. Only 7K scenes were used in this project as it was completed before the full release. is a large-scale dataset capturing diverse real-world scenes. We split it into $6723$ and $400$ scenes for training and performing evaluation, respectively. The $400$ testing scene is composed of the DL3DV benchmark (140 outdoor scenes) and 260 indoor scenes held out from its official training set to balance the indoor-outdoor ratio.
Hypersim [54] is a synthetic 3D indoor scene dataset with ultra photo-realistic renderings, aimed at testing the generalization capability to out-of-distribution indoor scenes. Hypersim is challenging due to its complicated geometry, extreme lighting conditions, and large camera baselines. We use a test set composed of 302 scenes.
MipNeRF360, Tanks & Temples (TT) [6, 36] includes 11 scenes for further testing the out-of-domain generalization capability of models on real data.
6.2 Evaluation and Baselines
We use $32$ views as input and use $32$ target novel view images for evaluation. The input and target views are non-overlap and are evenly sampled. We compare with three baseline methods:
GS-LRM and Long-LRM trained on DL3DV. These two baselines aim at validating the effectiveness of our proposed data. In detail, we train the GS-LRM and Long-LRM models on the largest real scene-level dataset, DL3DV, using the same training setting as ours.
Optimization-based 3DGS [35]. This baseline aims at validating the overall performance of our method, as the optimization-based 3DGS usually demonstrates a promising reconstruction quality. We use known camera information to get point cloud initialization from the $32$ input views using COLMAP. We use official training hyper-parameters.
Additionally, we note that comparing with more advanced 3DGS methods is not the focus our work, as our target is scaling up training data for improving feed-forward methods. 3DGS is only used to verify our overall performance, and our contributions can be directly ablated by comparing with LRMs trained without our data.
6.3 Results
Table 1 presents our results, demonstrating that training with both DL3DV and our MegaSynth dataset significantly improves model performance, with PSNR gains ranging from 1.2 to 1.8 dB. This improvement is consistent across model architectures (GS-LRM and Long-LRM), testing data (both in-domain DL3DV and out-of-domain datasets), image resolutions, and evaluation metrics, highlighting the effectiveness of our synthesized MegaSynth in enhancing the reconstruction quality of LRMs for wide-coverage scenes. Fig. 3 and Fig. 4 visually compare the reconstruction results for models with and without MegaSynth. We observe remarkable improvements in scenes with complicated scene structures, geoemtry, material and lighting, aligning with data generation designs (Sec. 4). Our approach also achieves substantially better results than the optimization-based 3DGS method while offering much faster inference speeds (e.g., over 2000 times speed-up from 5 minutes to 0.1 seconds).
We observe a notable trend of utilizing MegaSynth. The performance gains with MegaSynth on out-of-domain data are often larger than those on in-domain data. For example, Long-LRM achieves PSNR gains of 1.6 and 1.8 dB on Hypersim at resolutions of 128 and 256, respectively, surpassing the 1.3 and 1.4 dB improvements observed on the in-domain DL3DV dataset. GS-LRM results exhibits a similar pattern. The results underscore MegaSynth’s effectiveness in enhancing the generalization capability of LRMs.
6.4 Ablation Studies
In this section, we examine the impact of synthesized data quality, quantity, and properties on real-data performance, along with training paradigms for utilizing synthesized data. Without additional specification, the default experimental setup in analysis is the resolution-128 GS-LRM with pre-training + fine-tuning training protocol.
MegaSynth data quality, quantity and property. Table 2 presents our results. In general, we observe a positive correlation between performance of MegaSynth-only training and subsequent real-data fine-tuning, underscoring the value of MegaSynth in model training. Specifically, we refer MegaSynth-only training to the model trained after the per-training stage using only MegaSynth.
In Table 2 (0), training with a basic version of MegaSynth without controlling the data diversity and complexity results in lower performance than training with real data alone (Table 4), suggesting that unregulated synthesized data fails to enhance training. Additionally, we observe training instability, with pre-training and fine-tuning failing after around 70K iterations. We hypothesize that the high data randomness contributes to this instability, impeding effective learning and negatively affecting fine-tuning of real data.
Introducing control of data distribution, as shown in Table 2 (1), improves both MegaSynth-only training and real-data fine-tuning performance, emphasizing the importance of data quality and effectiveness of our data control method. However, training instability worsens, likely due to the increased complexity that amplifies training challenges.
Incorporating metadata during training mitigates this instability. Table 2 (2) shows that adding geometrical supervision, $\mathcal{L}_{\text{loc}}^{S}$, significantly improves stability and overall performance. This result underscores a key advantage of MegaSynth: the ability to provide additional ground-truth data. Expanding the dataset to include more scenes (i.e., 700K scenes in total), as in Table 2 (3), yields additional gains, showing the benefit of scale.
DL3DV-Test-Indoor DL3DV-Test-Outdoor
Data PSNR SSIM LPIPS PSNR SSIM LPIPS
DL3DV 25.41 0.853 0.150 23.09 0.771 0.183
DL3DV + MegaSynth 26.75 0.890 0.116 23.89 0.803 0.157
Table 3: Performance gains on indoor and outdoor test data. Results are from 128-resolution GS-LRM. The indoor test is our held-out set and the outdoor test is the DL3DV benchmark. Data details in Sec. 6.1.
Hypersim DL3DV-indoor
PSNR SSIM LPIPS PSNR SSIM LPIPS
Real-only 23.89 0.806 0.195 25.41 0.853 0.149
MegaSynth-only 21.50 0.719 0.272 23.02 0.782 0.217
Joint Training 25.33 0.844 0.157 26.54 0.885 0.121
Pre-training + Fine-tuning 25.46 0.846 0.154 26.75 0.890 0.116
Table 4: Ablation study on the training framework to leverage synthesized data. Results are reported on Hypersim and DL3DV indoor test subset with GS-LRM.
Table 5: Visual comparison between models trained with only MegaSynth and only real data. We include two failure cases of the MegaSynth-only model and highlight the failure regions. Results are reported with resolution 128 on Hypersim.
MegaSynth
PSNR SSIM LPIPS
w. Geomtry 23.24 0.684 0.280
w. Material 23.07 0.663 0.336
w. Lighting 22.65 0.660 0.305
Table 6: Analysis of the synthesized data control. Results are reported with resolution 128. The model is only trained on DL3DV to measure the distribution distance.
Table 7: Analysis of the NVS quality with different input view numbers. Results are reported with resolution 128 on Hypersim.
Indoor and outdoor improvements. We analyze the performance gains in Table 4, focusing on both indoor and outdoor test data from DL3DV. Although our synthesized MegaSynth data is primarily focusing on indoor scenes, we observe improvements in outdoor scenes as well, with a notably larger performance gain on indoor scenes. This suggests that the MegaSynth contributes to a generalized enhancement in geometric and appearance understanding, enabling broader generalization across diverse environments. At the same time, improving MegaSynth with outdoor characteristics would be an interesting direction.
Training strategies. We evaluate different strategies for utilizing MegaSynth. As shown in Table 4, training exclusively on MegaSynth (row 2) achieves performance comparable to training on real data (row 1), highlighting the effectiveness of MegaSynth and supporting our hypothesis that explicit semantics are not required for training scene reconstruction models. We visualize the results in Fig. 7. We find the model performs closely in most of the scenes but is much worse on complicated geometry patterns and large scene scales that are hard to model in synthesized data.
We further compare two approaches: (i) joint training on both synthesized MegaSynth and real data in row 3, and (ii) pre-training on MegaSynth followed by fine-tuning on real data in row 4. As shown in Table 4, the second approach yields slightly better performance, though the performance gap is minimal. This suggests that the model effectively learns the joint distribution of synthesized and real data without catastrophic forgetting during fine-tuning, indicating a degree of distribution alignment between MegaSynth and real data. Additionally, this experiment confirms that the performance gain results from the enhanced reconstruction capability acquired through MegaSynth, rather than simply from additional training iterations.
6.5 Analysis
We analyze the data control strategy and MegaSynth’s effectiveness with varying input views.
Analysis of data control. We evaluate the quality of MegaSynth and its alignment with real data by applying the GS-LRM model, trained on real-world DL3DV data, to our generated MegaSynth. As shown in Table 7, the reconstruction performance is reasonable with a 23 dB PSNR, indicating good alignment with real data. Notably, by controlling geometry, material, and lighting (introduced in Sec. 4) step by step, the performance slightly decreases, suggesting an increased data complexity.
Analysis on different numbers of input views. We extend our model trained with MegaSynth to scenarios with fewer input views, training GS-LRM with inputs of 8, 16, 24, and 32 views using either real-world data alone or a combination of real-world and MegaSynth data. As shown in Fig. 7, GS-LRM trained with both DL3DV and synthesized MegaSynth data demonstrates improved performance as the number of input views increases. Notably, an almost constant performance gap remains regardless of the number of views, which we attribute to the semantic gap between DL3DV and MegaSynth. These results highlight the effectiveness of MegaSynth for sparse-view reconstruction and suggest that semantic alignment is not a primary driver of 3D reconstruction performance.
7 Conclusion
We introduce MegaSynth, a non-semantic procedurally generated dataset, to improve LRMs for reconstructing wide-coverage scenes. MegaSynth benefits from its scalability and controllability, improving the model’s understanding of geometry and appearance. Experiments show MegaSynth’s capability of improving LRM reconstruction quality via both pre-training and joint training. The performance gains are consistent over different model architectures, test data domains, and input/output resolutions. Interestingly, LRMs trained sorely with MegaSynth demonstrate comparable performance with using real data, demonstrating that reconstruction is almost a non-semantic/low-level task.
Acknowledgements
The work was done while Hanwen Jiang, Desai Xie, Ziwen Chen, and Haian Jin were interns at Adobe Research. We thank Kalyan Sunkavalli for the support and feedback. Qixing Huang would like to acknowledge NSF IIS 2047677 and NSF IIS 2413161.
Appendix A MegaSynth Details
In this section, we include more details of our MegaSynth generation method. We introduce the details according to the sections in the main paper, i.e. scene floor plan, geometry and texture, and lighting.
A.1 Scene Floor Plan
We define the parameters of the scene size and object box in Table 8 and Table 9, including the categories, types, size ranges, height ranges, and probabilities. These object boxes are placed randomly in the scene, except for some categories, i.e. on-ground small box, on-roof box, and on-wall box, which have pre-defined location priors.
Scene parameters Size range [17.0, 30.0]
Height range [10.0, 15.0]
Object box parameters # Categories 7
Large object box Size range [4.0, 8.0]
Number range [2, 5]
Small object box Size range [2.0, 4.0]
Number range [4, 8]
Type 1 On-ground
Prob. 1 0.5
Height range 1 [2.0, 6.0]
Type 2 Atop large box
Prob. 2 0.5
Height range 2 [2.0, 4.0]
On-roof object box Size range [2.0, 5.0]
Number range [2, 4]
Type 1 Thin
Prob. 1 0.5
Height range 1 [0.5, 1.5]
Type 2 Thick
Prob. 2 0.5
Height range 2 [2, 4]
Table 8: Scene floor plan details part 1.
On-wall object box Size range [2.0, 5.0]
Number range [3, 6]
Type 1 Thin
Prob. 1 0.5
Height range 1 [0.5, 1.5]
Type 2 Thick
Prob. 2 0.5
Height range 2 [2, 4]
Wire-frame box Size range [3.0, 6.0]
Number range [1, 3]
Height range [3.0, 6.0]
Prob. 0.8
Thin stick box Length range [3.4, 18]
Type 1 On-wall
Prob. 1 1.0
Size 1 [0.1, 0.6]
Number 1 [5, 16]
Type 2 In-space
Prob. 2 0.5
Size 2 [0.8, 1.8]
Number 2 [2, 6]
Axis-aligned box Size range [2.0, 5.0]
Number range [1, 2]
Prob. 0.7
Height range [0.2, 1.0]
Table 9: Scene floor plan details part 2.
A.2 Geometry and Texture.
We include the details of object geometry in Table 10. In detail, the probability of using cube, sphere, cylinder and cone primitives are all 0.25 for large, small, on-wall, on-roof and wireframe object. For thin stick and axis-aligned objects, we only use cubes and cylinders. Beside, fro wireframe objects, we use cube, cylinder and torus, where torus has genus, increasing the geometry and topological complexity an diversity. We apply the height field augmentations to all shape primitives except for thin sticks and axis-aligned objects.
Large object Number of shape primitives [4, 5, 6, 7, 8, 9]
Prob. of Number of shape primitives [0.147, 0.206, 0.294, 0.206, 0.147]
Primitive types Default
Small object Number of shape primitives [2, 3, 4, 5]
Prob. of Number of shape primitives [0.25,0.375,0.25,0.125]
Primitive types Default
On-wall object Number of shape primitives [2, 3, 4, 5]
Prob. of Number of shape primitives [0.25,0.375,0.25,0.125]
Primitive types Default
On-roof object Number of shape primitives [2, 3, 4, 5]
Prob. of Number of shape primitives [0.25,0.375,0.25,0.125]
Primitive types Default
Wireframe object Number of shape primitives [1, 2, 3]
Prob. of Number of shape primitives [0.5, 0.25, 0.25]
Wireframe Primitive types Torus, Cube, Sphere
Wireframe thickness [mean scale/30, mean scale/20]
Sphere wireframes segments 8
Sphere wireframes ring count 8
Cube wireframes subdivision [1, 2, 3]
Cube wireframes subdivision prob. [0.33, 0.33, 0.33]
Torus wireframes minor radius 0.3 $\cdot$ mean scale
Torus wireframes major segments 8
Torus wireframes minor segments 8
ProB. Adding intersecting obj. 0.5
Types of intersecting obj. Default
Thin stick object Number of shape primitives 1
Primitive types Cube or Cylinder
Axis-aligned object Number of shape primitives 1
Primitive type Cube
Table 10: Object geometry details. ’mean scale’ is the average of the geometry size over the three axis.
We include the details of object textures in Table 11. After we randomly select textures and materials for all instantiated geometry primitives, we randomize the materials to improve complexity and diversity. We also have special deigns for materials of axis-aligned objects. We include details in Table 12.
Prob. modify mat. 0.5
Prob. modify mat. of slot 0.4
Prob. specular scene 0.2
Basic roughness range [0.001, 0.2]
Basic metallic range [0.001, 1.0]
Specular roughness range [0.0, 0.05]
Specular metallic range [0.6, 1.0]
Table 11: Material details 1.
Glass IOR range [1.4, 1.6]
Glass roughness range [0.001, 0.1]
Prob. Axis-aligned object glass 0.8
Table 12: Material details 2.
A.3 Lighting Details
We include the lighting details of sunlight in Table 13. We include details of luminous objects and light bulbs in Table 14.
Prob. sunlight 0.6
Sunlight strength [0.2, 2.0]
Prob. window glass 0.5
Prob. window bar 0.5
Table 13: Sunlight details.
Luminous objects Applied objects Thin sticks
Prob. 0.7
Prob. slot 0.2
Strength range 1 [0.2, 2.0]
Prob. strength range 1 0.9
Strength range 2 [5.0, 8.0]
Prob. strength range 2 0.1
Light bulb Num. range [2, 5]
Strength range 1 [0.2, 2.0]
Prob. strength range 1 0.9
Strength range 2 [5.0, 8.0]
Prob. strength range 2 0.1
Table 14: Luminous objects and light bulbs details.
Appendix B Model and Training Details
We include more model and training details as follows.
Training input and target view sampling. For each training sample in a batch, we randomly sample input views and target views from a pool of 48 views following LRM training strategy [27]. The number of input views is always 32. The number of target views is 12 for 128-resolution experiments, and 8 for 256-resolution experiments to balance the compute cost. We allow the overlap between input and target views during training. On the MegaSynth dataset, the set of 48 views are randomly sampled. On the real training data, we evenly sample frames within a distance range, which is sampled from the range of 64 to 128.
Camera pose normalization. The cameras of the input views are normalized with a random global scale between $1.1$ and $1.6$. For Gaussian rendering, we clip the predicted Gaussian scale of $0.135$. We set a near plane of the Gaussian renderer as $0.1$.
Learning rate and scheduler. In the pre-training stage, we use a peak learning rate of $44e-$ 4. In the tuning stage using real-world data, we use a smaller peak learning rate of $41e-$ 4. For joint training or training exclusively on real data, we use a learning rate of $44e-$ 4. All experiments adopt a warm-up of $3000$ iterations and cosine learning rate decay.
Batch size. For both $128\times 128$ and $256\times 256$ resolution training, we use a batch size of 4 per GPU. The experiments are launched on 64 A100 GPUs thus the global batch size is 256.
Training iterations. The training iterations for Res-128 and Res-256 are 120K and 80K for each training stage (i.e., pre-training and fine-tuning stages, as well as joint-training), respectively. The final learning rate is decreased to 0 at the end of training. Specifically, we end the pre-training stage at 75K and 55K iterations for experiments on resolution 128 and 256, respectively. Thus, The effective learning rate at the end of pre-training stage is around $41e-$ 4. The reason is we observe that training with more iterations, especially with a learning rate smaller than $41e-$ 4, leads to overfit on MegaSynth and makes the fine-tuning stage fail.
Training time cost. It takes GS-LRM $7$ days for pre-training and fine-tuning, and it takes $4$ days for joint-training, under resolution $128\times 128$. It takes $11$ days for pre-training and fine-tuning, and it takes $6$ days for joint training on resolution $256\times 256$.
Gaussian Settings. We use spherical harmonics of 3 for 3D Gaussians. We follow all other training hyper-parameters as the original GS-LRM [85] and Long-LRM [89].
Loss weights. We set the weights of point location loss (on synthetic data) and perceptual loss as 0.4 and 0.2, respectively. For joint training, we set the probability of sampling data from real and synthetic data as the same. For ablations, we run experiments with resolution $128\times 128$ using GS-LRM.
Training view rendering settings. For MegaSynth rendering, we sample 36 and 12 cameras in the outer and inner parts of the scenes, respectively. We sample the FoV of cameras within the range of 45 to 70 degrees.
Other details. In our ablation, quality control means we only use four basic object types without wireframes, think structure and axis-aligned object, without material randomization, and using only ambient lighting.
Appendix C More Results
We include more visualization results with 32 input views and 32 rendered target views as well as the ground-truth target views in Fig. 5 and Fig. 6.
Figure 5: Visualizaton of input views (first row of each example), render target view and ground-truth target views (last two rows of each example). We include results on the DL3DV benchmark data.
Figure 6: Visualizaton of input views (first row of each example), render target view and ground-truth target views (last two rows of each example). We include results on the Hypersim data.